OpenStudy (anonymous):

if I pick any single digit number and multiply it by 12345679. Then multiply the result by 9. How can I explain how this happens?

OpenStudy (dumbcow):

explain how what happens ??

OpenStudy (anonymous):

that is kind of cool, isn't it?

OpenStudy (anonymous):

its cool but but is there a trick to why this happens?

OpenStudy (aum):

123456789 x 9 = 1111111101 So if you pick any SINGLE digit number "a" and multiply it by 123456789 and then by 9 you are actually multiplying "a" by 1111111101 and you will get: aaaaaaaa0a

OpenStudy (anonymous):

one answer is because \(123456789\times 9=111,111,111\) but that might not be a very satisfying answer, just changes the question as to why that is so

OpenStudy (anonymous):

im still a little confused. No matter what single digit I multiply I will still get that number repeated 9 times..

OpenStudy (anonymous):

to see what the second thing is true, it is because we write out numbers in base ten what happens when you divide \(111,111,111\) by \(9\) ? try it and you will see why you get \(123,456,789\)

OpenStudy (anonymous):

@helpmath123 multiply \(123,456,789\) by \(9\) and see that you get \(111,111,111\) do you see it?

OpenStudy (anonymous):

Yea, I understand it a little more

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@dumbcow the result. Explaining what happens when I get the result

OpenStudy (aum):

Oh, I did not notice that 8 was omitted. So it is 12,345,679 * 9 = 111,111,111 a * 12,345,679 * 9 = a * 111,111,111 = aaa,aaa,aaa

OpenStudy (anonymous):

@aum oh now i get it. Thanks

OpenStudy (aum):

You are welcome.